OpenStudy (anonymous):

sqrt(3x^5)/sqrt(8x^2)

OpenStudy (anonymous):

hey, how's it going?

OpenStudy (anonymous):

good, have you helped me before?

OpenStudy (anonymous):

To see better \[\sqrt(3x^5)/\sqrt(8x^2)\]

OpenStudy (anonymous):

exactly!

OpenStudy (anonymous):

rewritten as \[\sqrt(3x^5/8x^2)\]

OpenStudy (anonymous):

simplified as \[\sqrt(3x^3/8)\]

OpenStudy (anonymous):

Rewritten as \[\sqrt(3x^3)/\sqrt(8)\]

OpenStudy (anonymous):

\[1/2*\sqrt(3/2)* \sqrt(x^3)\]

OpenStudy (anonymous):

where did the 1/2 come from?

OpenStudy (anonymous):

√3x3)/(√8) \[\sqrt(8)=2*\sqrt(2)\]

OpenStudy (anonymous):

oh ohkay.